OpenStudy (anonymous):

An insect crawls along the edge of a rectangular swimming pool of length 27m and width 21m. If it crawls from corner A to its opposite corner B which is half the total length of the pool in 30mins. A)what is its average speed? and B) What is the magnitude of its average velocity? I have the answers as A) 2.7cm/s and B) 1.9cm/s i have 4800cm/1800s = 2.7cm/s but i cannot figure out how they got 1.9cm/s for the average velocity can someone please explain this? Thanks

OpenStudy (anonymous):

Part A asks you to find its average speed, which is a scalar - it has no defined direction. In order to find the average velocity, it wants you to treat the sides of the pool as x and y components of a vector with direction pointing from A to B. 

OpenStudy (anonymous):

Sweet thanks i see it now. i have a^2*b^2=c^2=3400=cm/1800s=1.9cm/s Thanks again

OpenStudy (anonymous):

:)